# Pressure Drop in Non-Spherical Packed Beds: Influence of Geometry and Reynolds Number

### v2

@@ 4 Numerical Model
Each module was meshed separately, resulting in the grids for each geometry consisting of a number of separate regions, which were connected through a non-conformal coupling (NCC) mesh interface (greenshields2024), which joins the overlapping cells and imposes wall boundary conditions otherwise. Side walls of the module were treated as no-slip boundaries with standard zero Neumann boundary conditions for pressure. The top and bottom boundaries of each geometry were also treated using NCC interface, imposing periodic boundary conditions to replicate an infinitely repeating porous structure.
@@ 5.1 Validation
Figure 5: Comparison of normalized streamwise velocity fields $\overline{w}/\langle w\rangle$ obtained from PIV measurements and CFD simulations for two Reynolds numbers, $\mathrm{Re_{p}}=100$ and $200$. The upper figure (a) corresponds to P1, and the lower figure (b) corresponds to P3 within the packed bed module.
The numerical setup was validated by comparing results for $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ at $\mathrm{Re}_{\mathrm{p}}=100,200$ against experiments conducted by christin2025. We compared the mean flow fields at two planes located between the bars (P1 and P3), as shown in fig. 1. The experimental configuration (referred to as RC2) consists of 18 modules and an outlet section, with measurements available at layers 13th–17th and above the bed. Because the velocity at the 13th layer is least affected by outlet effects, it was selected for comparison with the fully developed, periodic flow of the simulation.
The flow field at P1 and P3 is compared in fig. 5, plotted in a $(\xi,z)$ coordinate system (horizontal and vertical coordinates, respectively). The simulation captures the experimental flow features with satisfactory accuracy, particularly at the midspan region. Both datasets indicate that the flow field behavior is mainly determined by the geometry of the void space, especially by the locations of the inlets and outlets. The accelerated fluid near these narrow connections between the modules generates several recirculation regions attached to the bottom and top walls.
@@
Figure 6: Normalized spanwise velocity profiles ($\overline{w}/\langle w\rangle$) from PIV measurements and CFD simulations. Profiles are shown at three vertical positions (z/B=12.25, 12.50, 12.75z/B=12.25,\,12.50,\,12.75) within the packed bed for $\mathrm{Re_{p}}=100$ and $200$.
@@ 5.2 Overview of the Flow Fields
For the rotation angle of $20\text{\,}\mathrm{\SIUnitSymbolDegree}$, visualized in fig. 7, the areas of high velocity field are confined to the central section of the void spaces. This is a direct result of a channel-like nature of the geometries with small rotation angle. Each void space is characterized by a strong streamwise connectivity (closely aligned and single inlet and outlet) inducing preferential transport of momentum along the centre. Moreover, a weak recirculating flow can be observed in the “cavity”like regions of the void spaces located at the sides of the strong core flow. This is consistent with the observations of Zhang2025, that in the inertial regime of laminar flow, recirculation zones emerge near the walls when the flow passes through pore-throat structures.
@@
After examining the influence of the rotation angle on the flow structure, the effect of the Reynolds number on the flow regime was considered. Figure 8 shows the normalized mean velocity magnitude for a geometry with $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ at two different Reynolds numbers. The slices shown in the fig. 8 are extracted from the mid-planes of three successive modules, located at $z=5$, $15$, and $25\text{\,}\mathrm{m}\mathrm{m}$, respectively.
At $\mathrm{Re}_{\mathrm{p}}=1$ (fig. 8), the flow is fully laminar and stationary with the pressure drop primarily governed by Darcy’s law. The fluid passes smoothly through the void spaces between layers, where the normalized velocity magnitude ranges from 0 to 2.5. The streamlines follow smooth slightly curved trajectories directed upward through the bed. The velocity distribution within the void space exhibits a low-speed, upward-directed flow that accelerates toward the center.
@@ 5.3 Friction Factor
In the viscous regime, all curves follow a hyperbolic trend with respect to $\mathrm{Re}_{\mathrm{p}}$, which appears as a straight line in the log–log scale. As the angle increases from $\alpha=$0\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ to $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the friction factor increases and reaches its maximum at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. Beyond this angle, the friction factor decreases gradually, and further changes in rotation angle no longer introduce a noticeable deviation. In other words, within this regime, all curves nearly collapse onto one another, indicating that the influence of the rotation angle becomes weak.
The specific alignment of the square bars at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ creates a severe constriction of the flow path. These narrowing void spaces increase the local shear stress, resulting in the maximum friction factor observed in the viscous regime.
@@
As the Reynolds number increases, deviations from the linear trend become apparent, signaling the onset of inertial effects and the transition toward the Forchheimer regime. Additionally, the angle yielding the maximum friction factor is found to shift with the Reynolds number. For instance, at $\mathrm{Re}_{\mathrm{p}}=200$, the maximum friction factor occurs at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, in contrast to the peak at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ observed in the viscous regime.
@@
At $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the flow shows behavior similar to contraction–expansion sequences in ducts. As the Reynolds number increases, the acceleration into void spaces creates local jets, leading to flow separation and recirculation downstream of these constrictions. Consequently, form drag becomes the dominant loss mechanism, resulting in the highest observed friction factor
One of the most notable trends is the nearly linear behavior for $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, where perfect orthogonality between modules allows the flow to develop with minimal obstruction, resulting in reduced flow resistance. Therefore, the $90\text{\,}\mathrm{\SIUnitSymbolDegree}$ configuration can also be classified as a channel-like geometry.
Figure 10 presents a comparison between the computed friction factor ($f_{\mathrm{p}}$), obtained from eq. 6, and the fitted non-dimensional friction factor ($f_{\mathrm{fitted}}$). The latter was evaluated using eq. 7, incorporating the calculated Darcy number and Forchheimer coefficient. As shown in fig. 10, the fitted curves closely follow the computed data over the selected range of Reynolds numbers. The agreement is particularly strong in the viscous regime, while slight deviations appear at higher Reynolds numbers, suggesting that additional simulations at higher $\mathrm{Re}_{\mathrm{p}}$ may be required to better capture the inertial effects.
5.4 Fitting Model Parameters
Figure 10: Comparison between the computed friction factor ($f_{\mathrm{p}}$) and the fitted friction factor ($f_{\mathrm{fitted}}$) as a function of $\mathrm{Re}_{\mathrm{p}}$ number for different angles. Each curve corresponds to a distinct angle ranging from $0\text{\,}\mathrm{\SIUnitSymbolDegree}$ to $90\text{\,}\mathrm{\SIUnitSymbolDegree}$, with the indicated scaling factors applied to separate the curves vertically for clarity.
Figure 11: Variation of (a) Darcy number and (b) Forchheimer coefficient w.r.t. the rotation angle ($\alpha$)
@@ 5.4 Fitting Model Parameters
The determination of the $\mathrm{Da}$ and $C_{\mathrm{F}}$ numbers was carried out in two steps using a least-squares fitting optimization procedure. First, the Darcy number was obtained for $\mathrm{Re}\leq 1$ by neglecting inertial effects (i.e. assuming $C_{\mathrm{F}}=0$) and using eq. 5. In the second step, the $C_{\mathrm{F}}$ was evaluated by fitting the eq. 5 over the entire Reynolds number range while keeping the Darcy number fixed.
The Darcy number shows a strong dependence on the rotation angle, demonstrating two distinct behaviours depending on the geometries as shown in fig. 11. For $\alpha\leq$20\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the structures are aligned in a channel-like form, which promotes a less perturbed flow and results in high permeability. Beyond this angle, the geometry becomes multiply connected, leading to a sharp drop in permeability. For $\alpha>$20\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the Darcy number exhibits a sinusoidal-like oscillation with increasing angle, showing reduced sensitivity to the rotation angle. Interestingly, although $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ produces a channel-like flow, its permeability does not follow the behaviour typically we observe in similar geometries. Instead, the Darcy number (or permeability) at $90\text{\,}\mathrm{\SIUnitSymbolDegree}$ remains significantly lower than that observed at small rotation angles such as $0\text{\,}\mathrm{\SIUnitSymbolDegree}$ or $5\text{\,}\mathrm{\SIUnitSymbolDegree}$. This discrepancy highlights the difficulty of predicting drag and permeability in such geometries.
@@
We also report the Forchheimer coefficient in fig. 11 as a function of the rotation angle. It should be noted that the inertial regime was not fully developed within the current Reynolds number range; higher Reynolds numbers would be required to accurately fit the inertial contribution of the pressure drop. Therefore, the presented correlations can be considered only valid within the stated range of Reynolds number. Figure 11 indicates that the highest Forchheimer coefficient, $C_{\mathrm{f}}$, occurs at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, which is consistent with the angle exhibiting the largest friction factor at $\mathrm{Re}_{\mathrm{p}}=200$.

### v3

@@ 4 Numerical model
Each module was meshed separately, resulting in the grids for each geometry consisting of a number of separate regions, which were connected through a non-conformal coupling (NCC) mesh interface (greenshields2024), which creates new set of faces at the intersections of overlapping faces on the coupled patches. These intersection faces become part of the finite volume mesh on both sides of the interface, ensuring that the mesh is topologically closed. As a consequence, conservation of mass and momentum is maintained by the standard finite volume discretisation, without any additional correction. The coupling is fully conservative and formally second-order accurate.
As with standard unstructured finite volume discretisation such boundary condition can still result in a decrease of the overall accuracy or convergence order, resulting from potentially high non-orthogonality and skewness of coupled faces, requiring detailed monitoring of the NCC regions. For some of the considered geometries, we have observed that the use of NCC conditioners led to the presence of pressure oscillations near the coupled boundaries. As their presence did not impede convergence and subsequent refinements near the coupled regions did not contribute to a significant change in averaged results (e.g., relative change of friction factor below 2%), we have deemed the accuracy offered by NCC acceptable.
Side walls of the module were treated as no-slip boundaries with standard zero Neumann boundary conditions for pressure. The top and bottom boundaries of each geometry were also treated using NCC interface, imposing periodic boundary conditions to replicate an infinitely repeating porous structure.
@@ 5.1 Validation
Figure 6: Comparison of normalised streamwise velocity fields $\overline{w}/\langle w\rangle$ obtained from PIV measurements and CFD simulations at $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ for $\mathrm{Re_{p}}=100$ and $200$. (a) Position P1, (b) position P3.
The numerical setup was validated by comparing results for $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ at $\mathrm{Re}_{\mathrm{p}}=100,200$ against experiments conducted by christin2025. We compared the mean flow fields at two planes located between the bars (P1 and P3), as shown in Fig. 1. The experimental configuration consists of 18 modules and an outlet section, with measurements available at layers 13th–17th and above the bed. Because the velocity at the 13th layer is least affected by outlet effects, it was selected for comparison with the fully developed, periodic flow of the simulation.
The flow field at P1 and P3 is compared in Fig. 6, plotted in a $(\xi,z)$ coordinate system (horizontal and vertical coordinates, respectively). The simulation captures the experimental flow features with satisfactory accuracy, particularly at the midspan region. Both datasets indicate that the flow field behaviour is mainly determined by the geometry of the void space, especially by the locations of the inlets and outlets. The accelerated fluid near these narrow connections between the modules generates several recirculation regions attached to the bottom and top walls.
@@
Figure 7: Normalised spanwise velocity profiles ($\overline{w}/\langle w\rangle$) from PIV measurements and CFD simulations at $\alpha=$30\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. Profiles are shown at three vertical positions (z/B=12.25, 12.50, 12.75z/B=12.25,\,12.50,\,12.75) for $\mathrm{Re_{p}}=100$ and $200$.
@@ 5.2 Overview of the flow fields
After examining the influence of the rotation angle on the flow structure, the effect of the Reynolds number on the flow regime was considered. Figure 9 shows the normalized mean velocity magnitude for a geometry with $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ at two different Reynolds numbers. The slices shown in Fig. 9 are extracted from the mid-planes of three successive modules, located at $z=5$, $15$, and $25\text{\,}\mathrm{m}\mathrm{m}$, respectively.
At $\mathrm{Re}_{\mathrm{p}}=1$ (Fig. 9), the flow is fully laminar and stationary with the pressure drop primarily governed by Darcy’s law. The fluid passes smoothly through the void spaces between layers, where the normalized velocity magnitude ranges from 0 to 2.5. The streamlines follow smooth, slightly curved trajectories directed upward through the bed. The velocity distribution within the void space exhibits a low-speed, upward-directed flow that accelerates toward the center.
@@ 5.3 Friction factor
Figure 11: Friction factor ($f_{\mathrm{eq}}$) as a function of $\mathrm{Re}_{\mathrm{eq}}$ for different rotation angles. The dotted black line shows the Ergun correlation (Eq. 11). Inset: same data computed with the single-bar diameter $d_{\mathrm{sb}}$.
@@
In the viscous regime, all curves follow a hyperbolic trend with respect to $\mathrm{Re}_{\mathrm{p}}$, which appears as a straight line in the log–log scale. As the angle increases from $\alpha=$0\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ to $25\text{\,}\mathrm{\SIUnitSymbolDegree}$, the friction factor increases and reaches its maximum at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. Beyond this angle, the friction factor decreases gradually, and further changes in rotation angle no longer introduce a large variation in $f_{\mathrm{p}}$.
@@
As the Reynolds number increases, deviations from the linear trend become apparent, signaling the onset of inertial effects and the transition toward the non-linear (Forchheimer) regime, for larger angles. However, the configuration at $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ exhibits a much slower transition to the inertial regime (see Section 5.6) with $f_{\mathrm{p}}$ resembling a straight line. Additionally, the angle yielding the maximum friction factor shifts with the Reynolds number. For instance, at $\mathrm{Re}_{\mathrm{p}}=200$, the maximum friction factor occurs at $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, in contrast to the peak at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ observed in the viscous regime.
@@
These values of the drag can be traced to the properties of each individual geometry. The specific alignment of the square bars at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ creates a severe constriction of the flow path. These narrowing void spaces increase the local shear stress, resulting in the maximum friction factor observed in the viscous regime. At $\alpha=$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the flow shows behaviour similar to contraction–expansion sequences in ducts. As the Reynolds number increases, the acceleration into void spaces creates local jets, leading to flow separation and recirculation downstream of these constrictions. Consequently, form drag becomes the dominant loss mechanism, resulting in the highest observed friction factor. In the case of $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the perfect orthogonality between modules allows the flow to develop with minimal obstruction, resulting in reduced flow resistance. Despite this low friction factor, the geometric parameters of the $90\text{\,}\mathrm{\SIUnitSymbolDegree}$ configuration ($D/d_{\mathrm{eq}}=3.23$, $a_{v}=0.193~\mathrm{mm^{-1}}$) place it within the lattice-like group rather than the channel-like regime observed at $\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$.
The accuracy of the Ergun depends strongly on the choice of characteristic length scale. Figure 11 presents the friction factor computed with the module-equivalent diameter $d_{\mathrm{eq}}$. In the viscous regime, the curves for the lattice-like geometries ($\alpha\geq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) collapse onto the Ergun correlation. This is attributed to the fact that $d_{\mathrm{eq}}$ incorporates the angle-dependent wetted surface area through $a_{v}$ (Eqs. 17 and 8), unlike $d_{\mathrm{sb}}$, which remains constant for all rotation angles. With increasing Reynolds number, the curves progressively diverge: at intermediate angles (notably $\alpha\approx$60\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) the friction factor exceeds the Ergun prediction, whereas the channel-like configurations ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) and $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ remain below it.
When the constant single-bar diameter $d_{\mathrm{sb}}$ is employed instead (inset of Fig. 11), the Ergun correlation overestimates the friction factor systematically across all angles and Reynolds numbers. Furthermore, the spread among the individual curves is considerably larger, since $d_{\mathrm{sb}}$ does not vary with $\alpha$ and is therefore unable to capture the differences in wetted surface area between configurations. These observations indicate that $d_{\mathrm{eq}}$ is the more suitable length scale for correlating the friction factor in the present geometry.
5.4 Determining permeability and forchheimer coefficient
@@
The determination of the $\mathrm{Da}$ and $C_{\mathrm{F}}$ numbers was carried out in two steps using a least-squares fitting optimization procedure of Eq. 7 to the gathered friction factor data. First, the Darcy number was obtained for $\mathrm{Re}\leq 1$ by neglecting inertial effects (i.e. assuming $C_{\mathrm{F}}=0$). In the second step, the $C_{\mathrm{F}}$ was evaluated by fitting the Eq. 7 over the entire Reynolds number range while keeping the previously determined Darcy number fixed.
Figure 12 presents a comparison between the computed friction factor ($f_{\mathrm{p}}$), obtained from Eq. 6, and the fitted non-dimensional friction factor ($f_{\mathrm{fitted}}$). The latter was evaluated using Eq. 7, incorporating the calculated Darcy number and Forchheimer coefficient. As shown in Fig. 12, the fitted curves closely follow the computed data over the selected range of Reynolds numbers. The agreement is strong in the viscous regime, while slight deviations appear at higher Reynolds numbers, suggesting that additional simulations at higher $\mathrm{Re}_{\mathrm{p}}$ may be required to better capture the inertial effects.
The computed Darcy number, plotted in Fig. 13, shows a strong dependence on the rotation angle, demonstrating two distinct behaviours depending on the geometry type. For $\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the gaps in consecutive modules are aligned in a channel-like form, which promotes a less perturbed flow and results in high permeability. As the angle increases, the connections between consecutive modules become more obstructed, leading to a sharp decrease in permeability with $\alpha$, observable also as a previously described trend in the friction factor in the viscous regime.
@@
The configurations at $\alpha=$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $20\text{\,}\mathrm{\SIUnitSymbolDegree}$ and $90\text{\,}\mathrm{\SIUnitSymbolDegree}$ are classified as lattice-like based on their bed-to-particle diameter ratio ($D/d_{\mathrm{eq}}>3$). This classification is confirmed by their Darcy numbers ($\mathrm{Da}=0.0020$, $0.0019$ and $0.0018$, respectively), which fall within the range of the lattice-like group ($\mathrm{Da}=0.0017\pm 0.0004$). For the $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ configuration in particular, $D/d_{\mathrm{eq}}=3.23$ and $a_{v}=0.193~\mathrm{mm^{-1}}$ are consistent with the lattice-like geometries, in contrast to the channel-like configurations at $\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, where $D/d_{\mathrm{eq}}<3$ and $a_{v}<0.16~\mathrm{mm^{-1}}$. This confirms that the bed-to-particle diameter ratio ($D/d_{\mathrm{eq}}\approx 3$ as boundary) and the hydraulic tortuosity ($\tau\approx 1.1$, see Fig. 14) provide a reliable basis for the classification.
Figure 12: Comparison between the computed friction factor ($f_{\mathrm{p}}$) and the fitted friction factor ($f_{\mathrm{fitted}}$) as a function of $\mathrm{Re}_{\mathrm{p}}$ for different angles. Scaling factors are applied to separate the curves vertically for clarity.
Figure 13: Darcy–Forchheimer model parameters as a function of the rotation angle $\alpha$. (a) Darcy number $\mathrm{Da}=K/B^{2}$, the tortuosity-based model ($K_{\tau}$) and Blake–Kozeny correlations using $d_{\mathrm{eq}}$ and $d_{\mathrm{sb}}$. The dashed line and shaded band indicate the DNS average ($0.0017$) and $\pm 2\sigma$ ($0.0004$), respectively. (b) Forchheimer coefficient $C_{\mathrm{F}}$ obtained from DNS fitting compared with values computed using $d_{\mathrm{sb}}$ and $d_{\mathrm{eq}}$.
The Forchheimer coefficient $C_{\mathrm{F}}$, shown in Fig. 13, exhibits a trend that parallels the Darcy number at small rotation angles: channel-like configurations ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) yield the lowest values of $C_{\mathrm{F}}$. This is not a direct consequence of the high permeability at these angles, but rather reflects a common geometric origin: aligned void spaces suppress flow separation and recirculation, thereby reducing both viscous resistance (high $\mathrm{Da}$) and inertial losses (low $C_{\mathrm{F}}$).
For $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, the low $C_{\mathrm{F}}$ is consistent with the delayed transition to the inertial regime observed for this configuration (see Section 5.6), which suggests that the Forchheimer term is not fully developed within the present Reynolds number range.
At intermediate angles ($30\text{\,}\mathrm{\SIUnitSymbolDegree}$$70\text{\,}\mathrm{\SIUnitSymbolDegree}$), $C_{\mathrm{F}}$ exhibits considerably stronger variations with $\alpha$ than at the extremes. In this range, the void-space topology changes significantly with each increment in rotation angle, producing markedly different levels of flow acceleration, separation and recirculation.
5.5 Modelling permeability and forchheimer coefficient
Figure 14: (a) $\tau_{\mathrm{eff}}$ (Eq. 24) as a function of $\mathrm{Re}_{\mathrm{p}}$ for different rotation angles. (b) Hydraulic tortuosity $\tau$ (Eq. 25) as a function of the rotation angle $\alpha$.
Figure 14(a) shows the upper bound for the hydraulic tortuosity $\tau_{\mathrm{eff}}$, computed from Eq. 24, as a function of the Reynolds number for all rotation angles. In the viscous regime, $\tau_{\mathrm{eff}}$ remains approximately constant for each angle, consistent with the Darcy flow assumption where the flow topology is independent of the flow rate. As the Reynolds number increases, inertial effects alter the flow paths and $\tau_{\mathrm{eff}}$ increases for most angles. The channel-like configurations ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$) exhibit the lowest values ($\tau_{\mathrm{eff}}\approx 1$), reflecting nearly straight flow paths. The $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ configuration also shows a low tortuosity ($\tau_{\mathrm{eff}}\approx 1.09$), placing it at the lower bound of the lattice-like group.
Figure 14(b) presents the hydraulic tortuosity $\tau$ (Eq. 25) as a function of the rotation angle. For $\alpha\leq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $\tau$ is close to unity, indicating that the aligned bar arrangement allows the flow to pass with minimal deviation. A sharp increase occurs between $\alpha=$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$ and $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, coinciding with the transition from channel-like to lattice-like geometries. For $\alpha\geq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, $\tau$ oscillates between approximately $1.1$ and $1.35$, with a peak near $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$. This is consistent with the friction factor results (Fig. 10), where the maximum drag in the viscous regime was also observed at $\alpha=$25\text{\,}\mathrm{\SIUnitSymbolDegree}$\$.
Figure 13 compares the DNS permeability with three model predictions. The Blake–Kozeny estimate based on the module-equivalent diameter, $K_{\mathrm{eq}}$ (Eq. 21), provides the closest agreement with the DNS data for the lattice-like geometries ($\alpha\geq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), with a mean absolute percentage error (MAPE) of $12.3\%$. Because $d_{\mathrm{eq}}$ incorporates the angle-dependent wetted surface area through $a_{v}$ (Eqs. 8 and 10), it captures the dominant geometric effect governing permeability variations in these configurations.
The tortuosity-based estimate $K_{\tau}$ (Eq. 23), which combines the hydraulic tortuosity $\tau$ with the specific surface area $a_{v}$ and a fitted Kozeny constant $c_{\mathrm{KC}}=3.1$, reproduces the angle-dependent trend with a MAPE of $18.9\%$. Both $K_{\mathrm{eq}}$ and $K_{\tau}$ provide good estimates of the permeability for the lattice-like geometries, with MAPEs below $20\%$. In contrast, $K_{\mathrm{sb}}$, based on the constant single-bar diameter, is nearly independent of $\alpha$ and yields a MAPE of $80.2\%$, confirming that it cannot reproduce the angle-dependent permeability.
For the channel-like geometries ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), all three models underestimate the permeability. These correlations are derived for multiply connected pore structures and do not account for the aligned, low-resistance void spaces present at small rotation angles.
The Ergun-based Forchheimer coefficients, $C_{\mathrm{F}}^{\mathrm{Erg}}(d_{\mathrm{sb}})$ and $C_{\mathrm{F}}^{\mathrm{Erg}}(d_{\mathrm{eq}})$, computed from Eq. 14 using $K_{\mathrm{sb}}$ and $K_{\mathrm{eq}}$ respectively, are compared with the DNS-fitted values in Fig. 13. Because Eq. 14 relates $C_{\mathrm{F}}$ directly to $\sqrt{K}$, the quality of the permeability estimate propagates into the predicted Forchheimer coefficient.
The single-bar estimate $C_{\mathrm{F}}^{\mathrm{Erg}}(d_{\mathrm{sb}})$ yields a constant value of approximately $0.7$, which underestimates the DNS-fitted $C_{\mathrm{F}}$ for the lattice-like geometries ($\alpha\geq$15\text{\,}\mathrm{\SIUnitSymbolDegree}$\$). Since $d_{\mathrm{sb}}$ does not vary with the rotation angle, the resulting $K_{\mathrm{sb}}$ cannot capture the angle-dependent permeability, and consequently the predicted Forchheimer coefficient inherits this limitation. The module-equivalent estimate $C_{\mathrm{F}}^{\mathrm{Erg}}(d_{\mathrm{eq}})$ varies with $\alpha$ through the angle-dependent specific surface area and reaches values closer to the DNS data at some intermediate rotation angles. However, it significantly overestimates $C_{\mathrm{F}}$ at small rotation angles ($\alpha\leq$10\text{\,}\mathrm{\SIUnitSymbolDegree}$\$), where the channel-like flow structure departs from the packed-bed assumptions underlying the Ergun correlation. At $\alpha=$90\text{\,}\mathrm{\SIUnitSymbolDegree}$\$, both Ergun estimates also overestimate $C_{\mathrm{F}}$, consistent with the low DNS-fitted value discussed above, which is attributed to the delayed transition to the inertial regime at this angle. Moreover, the DNS data are limited to $\mathrm{Re}_{\mathrm{p}}\leq 200$, which may not be sufficient to fully establish the Forchheimer coefficient for all geometries. The development of geometry-aware models for the Forchheimer coefficient, supported by simulations at higher Reynolds numbers, remains a subject for future investigation.